a) Solve the differential equation
dy/dx+xy=y^2e^(x^2/2)sinx
1
Expert's answer
2020-03-24T13:44:00-0400

"y'+xy=y^2e^{x^2/2}\sin{x}"

divide both sides by "y^2"

"y'/y^2+x/y=e^{x^2/2}\sin{x}"

substitution: "t=1\/y"

"xt-t'=e^{x^2/2}\sin{x}"

let's solve first this equation: "xt-t'=0"

"\int dt/t=\int xdx"

"\ln{t}-\ln{C}=x^2/2"

"t=Ce^{x^2\/2}"

then solution to initial equation will have form: "t(x)=C(x)e^{x^2\/2}"

"t'(x)=C'(x)e^{x^2\/2}+xC(x)e^{x^2\/2}"

Let's put these to initial equation

"xC(x)e^{x^2/2}-C'(x)e^{x^2/2}-xC(x)e^{x^2/2}=e^{x^2/2}\sin{x}"

"C'(x)=-\sin{x}"

"C(x)=\int -\sin{x}dx=C_1+\cos{x}"

then "t(x)=(C_1+\cos{x})e^{x^2/2}"

"y=1/t=e^{-x^2/2}/(C_1+\cos{x})"